OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) − 10 = 12 + 4(x − 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

too late, too much work, my mind wont work :P sorry

OpenStudy (anonymous):

lol I can help

OpenStudy (anonymous):

You first have to distribute.

OpenStudy (anonymous):

2(x + 6) − 10 = 12 + 4(x − 1) 2x + 12 - 10 = 12 + 4x -4 -10 - 10 -4 -4 2x + 2 = 8 + 4x

OpenStudy (anonymous):

distribute so , 2x+12-10 = 12 +4x-4, combine like terms 2x+2 = 8+4x minus the 2x over and minus the 8 over -6 = 2x divide by 2 and you get x = -3

OpenStudy (anonymous):

2x + 2 = 8 + 4x Now combine like terms. -2x -2x ---------------- 2 = 8 + 2x -8 -8 ------------- -6 = 2x /2 /2 ------------------- x = -3